WARNING: This web is not used anymore. Please use PDBService.MinuteS12January06 instead!
 

Minutes 12 January 06

  • Phone: Lee(CMS), Stefan(ATLAS), Richard(ATLAS), Gancho(ATLAS), Iris(BNL), Gordon(RAL), Doris(GridKA), Silke(GridKA), Yuri(BNL), Sasha(ATLAS)
  • @CERN: Vicenzo(CMS), Zhen(CMS), Marco Clemencic(LHCb), Dirk(3D), Luca(CERN DB), Jacek(CERN DB), Luis(3D), Giacomo(POOL), Eva(3D)
  • Apologies: Barbara(CNAF)

  • General points:
    • Weekly meeting confirmed
    • DB readiness workshop 6 Feb (Mon) at CERN (1 day)
      • US talks will be during the afternoon
      • Dirk will send a reminder
    • Next DB workshop (DBA meeting) after CHEP
      • Volunteers: Gordon offered to host it at RAL
      • Dirk/Gordon will look at suitable dates
    • There will be a presentation about Frontier Caches next week (Luis)
      • Lee will provide input on CMS application rules/conventions to insure consistency
  • Frontier/Squid setup at Tier0
    • Dirk asked for the machines: 3 boxes will arrive during next weeks
    • GigabitEthernet network
    • 3D machine had problems during Xmas. Hardware intervention schedules for tomorrow
    • 2 disk servers for the production phase: installation before workshop
  • Rac and Storage setup at Tier0 (Luca and Jaceck)
    • Slides available on the agenda pages
    • More documentation about installation and post tasks available on wiki pages (see link on the slides)
    • Questions: ask Luca by email and use 3D mailing lists to keep logs about the discussions
  • 10gR2 upgrade for the testbed to be scheduled
  • POOL/Frontier performance in the testbed (Lee)
    • Slides available on the agenda pages
  • Frontier plugin releases coming up during the next 2 weeks
  • Roundtable Tier1 sites status
    • BNL (Iris): Oracle Cluster 10gR2 installed, ASMLib 2.0. Performance tests ongoing
    • RAL (Gordon): hardware ordered but not arrived. Intallation to February
    • CNAF (Barbara): Cluster setup done, OCFS but plan to go to ASM and Enterprise Server
    • GridKA (Doris): Expect production server hardware in February - March. Number of nodes as in the wiki table
    • No progress on the network problem with GridKA
    • IN2P3 is working correctly now. No streams backlog - so ARDA latency tests can continue
  • Roundtable testbed test progress
    • Stefan: just back at work - will progress with the Cool tests at RAL now
    • Marco: connected to RAL but no further tests yet
    • ATLAS online tests to be finished by Dirk. Meeting with Richard and Gancho setup to discuss workplan
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2006-01-13 - DirkDuellmann
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback